### v3

# FLIP: Towards Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

Abstract
Click-through rate (CTR) prediction plays as a core function module in various personalized online services. The traditional ID-based models for CTR prediction take as inputs the one-hot encoded ID features of tabular modality, which capture the collaborative signals via feature interaction modeling. But the one-hot encoding discards the semantic information conceived in the original feature texts. Recently, the emergence of Pretrained Language Models (PLMs) has given rise to another paradigm, which takes as inputs the sentences of textual modality obtained by hard prompt templates and adopts PLMs to extract the semantic knowledge. However, PLMs generally tokenize the input text data into subword tokens and ignore field-wise collaborative signals. Therefore, these two lines of research focus on different characteristics of the same input data (i.e., textual and tabular modalities), forming a distinct complementary relationship with each other. In this paper, we propose to conduct Fine-grained feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. We design a novel joint reconstruction pretraining task for both masked language and tabular modeling. Specifically, the masked data of one modality (i.e., tokens or features) has to be recovered with the help of the other modality, which establishes the feature-level interaction and alignment via sufficient mutual information extraction between dual modalities. Moreover, we propose to jointly finetune the ID-based model and PLM for downstream CTR prediction tasks, thus achieving superior performance by combining the advantages of both models. Extensive experiments on three real-world datasets demonstrate that FLIP outperforms SOTA baselines, and is highly compatible for various ID-based models and PLMs. The code is available [^1] https://github.com/justarter/FLIP.
Index Terms:
Fine-grained Alignment, Pretrained Language Model, CTR prediction, Recommender Systems
I Introduction
Click-through rate (CTR) prediction is the core component of various personalized online services (e.g., web search [1, 2, 3], recommender systems [4, 5]). It aims to estimate a user’s click probability towards each target item, given a particular context [6, 7, 8]. Recently, the emergence of Pretrained Language Models (PLMs) [9] has facilitated the acquisition of extensive knowledge and enhanced reasoning abilities, hence introducing a novel paradigm for CTR prediction task [10].
The traditional ID-based models for CTR prediction adopt the one-hot encoding to convert the input data into ID features, which we refer to as tabular data modality. An example is shown as follows:
$\underbrace{[0,0,1,\cdots,0]}_{\texttt{UserID=02}}\ \underbrace{[0,1]}_{\texttt{Gender=Male}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{ItemID=01}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{Genre=Action}}$ (1)
Deriving from POLY2 [11] and FM [12], the key idea of these models is to capture the high-order feature interaction patterns across multiple fields with different operators (e.g., product [13, 14], convolution [15, 16], and attention [17, 18]). On the other hand, the development of PLMs has given rise to another model paradigm that utilizes the PLM as a text encoder or recommender directly. The input data is first transformed into textual sentences via hard prompt templates, which could be referred to as textual data modality. In this way, the semantic information among original data is well preserved as natural languages (e.g., feature as text “male” instead of ID code “[0,1]”). Then PLMs [19, 20] encode and understand the textual sentences, and thus turn CTR prediction into either a binary text classification problem [21, 22] or a sequence-to-sequence task [23, 24] for inferring user preference. However, both of the above models possess inherent limitations.
In terms of ID-based models, they utilize the one-hot encoding that discards the semantic information conceived in the original features and fails to capture the semantic correlation between feature descriptions [25, 26]. In addition, ID-based models rely heavily on the user interactions and may not work well in scenarios with sparse interactions [27]. The aforementioned issues can be effectively alleviated by PLMs. PLMs have the ability to extract personalized information from feature descriptions, and use their external knowledge and reasoning capabilities to achieve robust performance in sparse-interaction situations [28, 29].
Furthermore, PLMs also have certain limitations. Since the input data is formulated as textual sentences which would be broken down into subword tokens, language models generally lose sight of field-wise collaborative signals among multiple features [30, 31], which are indispensable in CTR prediction. Additionally, it is hard to distinguish features that exhibits little textual variations in natural language [32, 33] (e.g., in terms of the movie name, “The Room” and “Room” are two very similar movies literally), which hinders language models from capturing subtle similarities and differences between different feature descriptions [34, 35]. Fortunately, the ID-based model can distinguish each different feature and perceive field-wise collaborative signals with the distinctive encoding method and model structure.
Figure 1: Three cross-modal pretraining tasks. Task (a) provides coarse-grained instance-level alignment via contrastive learning, while tasks (b) and (c) achieve fine-grained feature-level alignment through joint reconstruction.
As discussed above, both types of models possess their own advantages and disadvantages. So it is evident to construct a hybrid model that aligns and exploits both collaborative and semantic signals from ID-based models and PLMs, respectively. Meanwhile, in order to effectively solve the above challenges, the alignment of the two modalities needs to be refined to the feature level, so that ID-based models can perceive semantic information corresponding to each feature and PLMs can clearly distinguish features with similar text but different IDs.
However, as shown in Figure 1(a), most existing methods [25, 30] perform the contrastive learning [34] to align the ID-based models and PLMs. The Instance-level Contrastive Learning (ICL) is only coarse-grained and are not sufficient to solve the above feature-level challenges. In addition, ICL can not sufficiently capture the underlying relationships between features. It results in the representation degeneration problem [36, 37, 38], which will degrade the performance. We illustrate the feature embeddings learned by ICL in Figure 2(a), which collapse into a narrow cone and lead to limited expressive capacity. In contrast, our proposed FLIP is able to learn more distributed representations, and therefore better capture the feature diversity as shown in Figure 2(b).
To this end, in this paper, we propose to conduct Fine-grained feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. FLIP is a model-agnostic framework that adopts the common pretrain-finetune scheme [7, 19]. As illustrated in Figure 1(b) and 1(c), for pretraining objectives, we propose to build joint masked tabular and language modeling, where the masked features (i.e., ID or text of some features) of one modality are recovered with the help of another modality. This is motivated by the fact that both tabular and textual data convey almost the same information of the original raw data, but only in different formats. In order to accomplish the masked feature reconstruction, each single model (ID-based model or PLM) is required to seek and exploit the valuable knowledge embedded in the other model that corresponds to the masked features. In this way, the fine-grained feature-level alignment between feature IDs and texts is established.
(a) ICL
(b) Our FLIP
Figure 2: Visualization of feature embeddings on MovieLens-1M dataset. We use SVD [39] to project the feature embedding matrix into 2D data. (a) Instance-level Contrastive Learning (ICL) learns feature embeddings that collapse into a narrow cone, resulting in limited expressive capacity. (b) Our FLIP learns feature embeddings that are more distributed, therefore better capturing the feature diversity.
Then, as for finetuning, the fine-grained alignment pretraining has provided an enhanced modality-aware parameter initialization. Hence, we propose to jointly tune ID-based model and PLM to achieve superior performance. In the experiment, we further explore the effectiveness of finetuning only one of two models.
The main contributions of this paper are as follows:
  • We are the first to highlight the importance of utilizing fine-grained feature-level alignment between ID-based models and PLMs, in order to explore the potential of enhancing the recommendation performance of existing recommender systems.
  • We propose a model-agnostic framework called FLIP, in which the joint masked tabular and language modeling tasks are involved. In these tasks, the masked features of one modality have to be recovered with the help of another modality, thus establishing feature-level interactions and alignments between two modalities.
  • Extensive experiments on three real-world public datasets demonstrate the superiority of FLIP, compared with existing baseline models. Moreover, we validate the model compatibility of FLIP in terms of both ID-based models and PLMs.
II Related Work
In this section, we will review previous works which are highly related to ours. We will mainly focus on two aspects, including ID-based models and PLMs for CTR prediction.
II-A ID-based Models for CTR Prediction
Click-through rate (CTR) prediction serves as a core function module in personalized online services, including online advertising, recommender systems, etc [40]. Traditional ID-based CTR models follow the common design paradigm: embedding layer, feature interaction (FI) layer, and prediction layer. These models take as input one-hot features of tabular data modality, and employ various interaction functions to capture collaborative signals among features.
Earlier methods are limited in their ability to model fixed-order or low-order feature interactions. For instance, FM-based approaches [41, 42] employ decomposition parameters to model pairwise interactions. Due to the significance of FI in CTR prediction, numerous studies focus on designing novel structures for the FI layer to capture more informative and complex feature interactions. Wide&Deep [43] combines a wide network and a deep network to achieve the advantages of both. DCN [44] and DCNv2 [14] improve Wide&Deep by replacing the wide part with a more expressive cross network to learn explicit high-order feature interactions. DeepFM [5] combines DNN and FM, and xDeepFM [45] extends DeepFM by using a compressed interaction network (CIN) to capture feature interactions in a vector-wise way. Furthermore, various explicit or implicit interaction operators are designed to improve performance. For instance, PNN [13] uses the productive operator, AFN [46] utilizes the logarithmic operator, AFM [18] and AutoInt [17] use the attention operator.
Although the works above can capture the collaborative knowledge effectively, they disregard the original semantic information of the features and rely heavily on extensive user interactions, which constrains their potential applications and hinders their performance.
II-B PLMs for CTR Prediction
Pretrained Language Models (PLMs), as an emerging technology in the field of Natural Language Processing (NLP), are equipped with a vast number of parameters and undergo their rigorous pretraining process [19, 20]. PLMs demonstrate exceptional success in a wide range of tasks, owing to their extensive knowledge and strong reasoning abilities [47, 48, 49]. Inspired by these achievements, the application of language models for recommender systems has received more and more attention and research [10, 50, 51, 52, 53, 54, 55, 56, 57]. Different from one-hot encoding in ID-based models, this line of research needs to convert the raw data into textual modality, thus retaining the original semantic information of features.
Recent efforts to adapt PLMs for CTR prediction have yielded several breakthroughs. For instance, CTR-BERT [22] leverages a two-tower structure, and feeds user and item information into their respective BERT [19] towers for final CTR prediction. PTab [21] pretrains a BERT model based on masked language modeling using dataset text, and then finetunes this BERT on downstream discriminative tasks. P5 [23] uniformly converts different recommendation tasks into text generation tasks where the backbone model is T5 [58]. M6-Rec [59] utilizes M6 [60] as the backbone for downstream adaptation by representing user behavior data as plain texts and converting the tasks to language generation. Recformer [26] utilizes a bidirectional Longformer [61] to encode items which are formulated as textual key-value attribute pairs. Also, TALLRec [31] and related work [62, 63] utilize decoder-only large language models (e.g., LLaMA [64]) to follow instructions and answer the user preference question that are added after the textual input sentences. Other works also explore the application of language models in different settings and scenarios (e.g., zero-shot [65], few-shot [66], fairness [67], and graph learning [29]).
PLMs can extract personalized text information and harness their knowledge and reasoning capabilities to inform ID-based models, thereby amplifying their recommendation performance. However, it is challenging for PLMs to explicitly capture the field-wise collaborative signals among features and distinguish features with subtle textual differences, resulting in a non-negligible performance decline.
To exploit the advantages of both ID-based models and PLMs, existing state-of-the-art methods, such as CTRL [25] and RLMRec [30], utilize the instance-level contrastive learning (ICL) to align these two models. However, the ICL task is coarse-grained and insufficient to promote the model capacity. It also fails to adequately capture the underlying patterns and relationships among features, resulting in representation degeneration problems.
Therefore, we propose to conduct fine-grained feature-level alignment between ID-based models and PLMs via the joint masked tabular and language modeling pretraining tasks. In the finetuning stage, we propose to tune both models, and thus make full use of the advantages of both textual and tabular data modalities to achieve superior CTR prediction performance.
III Preliminaries
III-A ID-based Models for CTR Prediction
In tradition CTR prediction settings, we model the problem as a binary classification task [5, 13]. The dataset with $N$ instances is presented as $\{(\textbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{0,1\}$ (click or not) is the true label indicating user’s actual click behavior. The input $\textbf{x}_{i}$ is usually categorical tabular data with $F$ different fields, which can be formulated as $\textbf{x}_{i}=[x_{i,1},x_{i,2},\ldots,x_{i,F}]$. The goal of CTR prediction is to estimate the click probability $P(y_{i}=1|\textbf{x}_{i})$ based on input $\textbf{x}_{i}$. Generally, traditional ID-based models adopt the common “Embedding & Feature Interaction” paradigm [17, 45], which can be formulated as the following three layers.
Embedding Layer. The input $\textbf{x}_{i}$ is first transformed into high-dimensional sparse features via one-hot encoding [68]. Then we apply the embedding layer to map the one-hot vectors to dense low-dimensional embedding vectors $\textbf{E}_{i}=[\textbf{e}_{i,1},\textbf{e}_{i,2},\cdots,\textbf{e}_{i,F}]\in\mathbb{R}^{F\times D}$, where $D$ is the embedding size, and each feature $x_{i,f}$ is represented as a fixed-length vector $\textbf{e}_{i,f}\in\mathbb{R}^{D}$.
Feature Interaction (FI) Layer. The FI layer takes as inputs the embedding vectors E, and generates a compact representation $\textbf{v}_{i}$ for each instance $\textbf{x}_{i}$. Different structures are designed to capture the feature interaction patterns with different operators (e.g., MLP [5, 43], product [13, 44, 14], and attention [17, 69]).
Prediction Layer. The prediction layer (usually a linear layer or an MLP module) estimates the click probability $\hat{y_{i}}\in[0,1]$ based on the compact representation $\textbf{v}_{i}$, where $\sigma(x)=1/(1+\text{exp}(-x))$ is the sigmoid function. Finally, the model is trained with the binary cross-entropy (BCE) loss in an end-to-end manner:
$\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum\nolimits_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right]$ (2)
III-B PLMs for CTR Prediction
As PLMs have shown remarkable success in a wide range of tasks due to their extensive knowledge and reasoning capabilities [20, 49], researchers now tend to leverage their proficiency in natural language understanding and world knowledge modeling to solve the CTR prediction task [10, 50].
Different from ID-based models, (1) PLMs first transform the input data $\textbf{x}_{i}$ into the textual sentence $\textbf{x}^{text}_{i}$ via hard prompt templates. (2) Then the textual sentence $\textbf{x}^{text}_{i}$ is converted into meaningful lexical tokens through tokenization, which are embedded into a vector space. (3) Next, PLMs utilize their transformer-based neural networks to process the token embeddings obtained from the previous step, generating contextually coherent representation $\textbf{w}_{i}$. (4) Finally, we can either add a randomly initialized classification head (usually MLP) on the representation $\textbf{w}_{i}$ to perform binary classification and predict the click label $y_{i}\in\{0,1\}$ [21, 25], or add a language modeling head to do causal language modeling tasks and predict the likelihood of generating the next keyword (e.g., “yes” or “no”) through a verbalizer [23, 31].
IV Methodology
In this section, we present the FLIP framework, where fine-grained feature-level alignments between PLMs and ID-based models are established to promote the downstream CTR performance.
Figure 3: An illustrative example of modality transformation, which converts tabular data into textual data. We first construct a descriptive sentence for each feature field, and then concatenate them to obtain the final textual data.
IV-A Overview of FLIP
Figure 4: The overall framework of our proposed modality alignment pretraining. The instance-level contrastive learning (ICL) module provides coarse-grained instance-level self-supervisions. The joint reconstruction of masked language/tabular modeling (MLM/MTM) requires to recover the masked feature information with the help of the other modality, thereby enabling fine-grained feature-level alignments.
Our proposed FLIP framework follows the pretrain-finetune scheme, which can be divided into the following three major stages:
  • Modality transformation, as depicted in Figure 3, converts the raw data of tabular modality into textual modality. Note that inputs $\textbf{x}_{i}^{tab}$ and $\textbf{x}_{i}^{text}$ contain almost the same information of raw data, but only in different modalities.
  • Modality alignment pretraining, as shown in Figure 4, explicitly interacts between the dual modalities. Specifically, while the instance-level contrastive learning (ICL) module provides coarse-grained instance-level self-supervisions, the joint reconstruction of masked language/tabular modeling (MLM/MTM) could lead to fine-grained feature-level alignments between two models.
  • Supervised finetuning, as illustrated in Figure 5, further adapts the aligned PLM and ID-based model to the downstream task for final CTR prediction.
Hereinafter, we omit the detailed structure of PLMs and ID-based models, since FLIP serves as a model-agnostic framework and is compatible for various backbone models.
IV-B Modality Transformation
Standard PLMs take sequences of words as inputs [19, 47]. Consequently, the modality transformation aims to convert the tabular data $\textbf{x}_{i}^{tab}$ into textual data $\textbf{x}_{i}^{text}$ via hard prompt templates. Previous works [25, 21, 70] have suggested that sophisticated templates (e.g., with more vivid descriptions) for textual data construction might mislead the language model and make it fails to grasp the key information in the texts. Hence, we adopt the following simple yet effective transformation template:
$t_{i,f}$ $=[m_{f},\text{``}is\text{''},v_{i,f},\text{``}.\text{''}],\;f\in\{1,\cdots,F\}$ (3)
$\textbf{x}^{text}_{i}$ $=[t_{i,1},t_{i,2},\cdots,t_{i,F}],$
where $m_{f}$ is the name of $f$-th field (e.g., gender), $v_{i,f}$ denotes the feature value of $f$-th field for input $x_{i}^{tab}$ (e.g., female), and []delimited-[][$\cdot$][ ⋅ ] indicates the concatenation operator.
An illustrative example is given in Figure 3, where we first construct a descriptive sentence $t_{i,f}$ for each feature field, and then concatenate them to obtain the final textual data $\textbf{x}_{i}^{text}$. In this way, we preserve the semantic knowledge of both field names and feature values with minimal preprocessing and no information loss.
IV-C Modality Alignment Pretraining
After obtaining the paired text-tabular data $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$ for the same original raw input, we design a novel modality alignment pretraining algorithm to achieve both instance-level and feature-level alignments for PLMs and ID-based models.
As shown in Figure 4, we first perform field-level data masking to obtain the corrupted version of input pair, i.e., $(\hat{\textbf{x}}_{i}^{text},\hat{\textbf{x}}_{i}^{tab})$. Then, the PLM $h_{\text{PLM}}$ and ID-based model $h_{\text{ID}}$ encode the input data to obtain the dense representations $(\textbf{w}_{i},\hat{\textbf{w}}_{i})$ and $(\textbf{v}_{i},\hat{\textbf{v}}_{i})$ for textual and tabular modalities, respectively. Next, we apply three different pretraining objectives over the dense representations for loss computation and model optimization:
  • Joint reconstruction for Masked Language Modeling (MLM) requires the model to recover the original tokens from the corrupted textual context with the help of complete tabular data.
  • Joint reconstruction for Masked Tabular Modeling (MTM) requires the model to recover the original feature IDs from the corrupted tabular context with the help of complete textual data.
  • Instance-level Contrastive Learning (ICL) draws positive samples together (i.e., different modalities of the same input) and pushes apart negative sample pairs.
While ICL only provides coarse-grained self-supervisions, the joint reconstruction for MLM/MTM casts a dual pretraining task with fine-grained self-supervised signals. That is, each single model (i.e., PLM or ID-based model) has to seek and exploit the useful knowledge underlying the representation from the other modality to accomplish the recovery task, resulting in fine-grained feature-level cross-modal alignments.
IV-C1 Field-level Data Masking
We introduce the field-level data masking strategy for textual and tabular data, respectively.
For textual data, we propose to mask tokens at the field level. We first randomly select a certain ratio $r_{text}$ of the feature fields to be corrupted, and then mask all the consecutive tokens that constitute the entire text of the corresponding feature value. Note that this is quite different from the original random token-level masking strategy for common language models (e.g., BERT [19]), which is not a good choice for FLIP. The hard template we adopt has fixed patterns and structures, i.e., $\textbf{x}_{i}^{text}$ is simply a series of “what is what” sentences, and the descriptive text for each feature is also fixed. Hence, a random token-level masking strategy might lead to a too-easy pretraining task, where the model could suffer from the early degeneration problem and lose the generalization ability [71]. Besides, we denote the index set of masked tokens as $\mathcal{I}^{text}$.
As an example for textual data masking, suppose the tokenization of sentence for occupation field is [“occupation”, “is”, “college”, “student”]. The outcome of field-level masking should be [“occupation”, “is”, [MASK], [MASK]]. If the token-level masking is adopted, the outcome might be [[MASK], “is”, “college”, “student”] or [“occupation”, “is”, “college”, [MASK]]. Obviously, these masked tokens can be easily inferred solely based on the textual context, thus limiting the cross-modal interaction for joint reconstruction.
For tabular data, following previous works [7, 72], we uniformly sample a certain ratio $r_{tab}$ of the fields, and then replace the corresponding feature with an additional <MASK> feature, which is set as a special feature in the embedding table of ID-based model. The <MASK> feature is not field-specific, but shared by all feature fields to prevent introducing prior knowledge about the masked field [7]. Similarly, we denote the index set of masked fields as $\mathcal{I}^{tab}$.
After the field-level data masking, we obtain the corrupted version for both modalities, i.e., $(\hat{\textbf{x}}_{i}^{text},\hat{\textbf{x}}_{i}^{tab})$. It is worth noting that the selected fields to be masked do not necessarily have to be the same for textual and tabular modalities, since they serve as two independent parallel workflows.
IV-C2 Data Encoding
We employ the PLM $h_{\text{PLM}}$ and ID-based model $h_{\text{ID}}$ to encode the input pairs from textual and tabular modalities, respectively:
$\textbf{w}_{i}$ $=h_{\text{PLM}}(\textbf{x}_{i}^{text}),\;\;\;\hat{\textbf{w}}_{i}=h_{\text{PLM}}(\hat{\textbf{x}}_{i}^{text}),$ (4)
$\textbf{v}_{i}$ $=h_{\text{ID}}(\textbf{x}_{i}^{tab}),\;\;\;\;\;\;\;\hat{\textbf{v}}_{i}=h_{\text{ID}}(\hat{\textbf{x}}_{i}^{tab}).$
Here, $\textbf{w}_{i}=[w_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ is the set of hidden states from the last layer of language model, where $L$ is the number of tokens for $\textbf{x}_{i}^{text}$ and $D_{text}$ is the hidden size of language model. Note that $w_{i,1}$ is the [CLS] token vector that represents the overall textual input. $\hat{\textbf{w}}_{i}=[\hat{w}_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ also satisfies the notations above. $\textbf{v}_{i},\hat{\textbf{v}}_{i}\in\mathbb{R}^{D_{tab}}$ are the representations produced by ID-based model.
IV-C3 Joint Reconstruction for Masked Language Modeling
As shown in Figure 4, the MLM module takes the text-tabular pair $(\hat{\textbf{w}}_{i},\textbf{v}_{i})$ as input, and attempts to recover the masked tokens. We denote the index set of masked tokens as $\mathcal{I}^{text}$. For each masked token with index $l\in\mathcal{I}^{text}$, we concatenate the corresponding token vector $\hat{w}_{i,l}$ with the “reference answer” $\textbf{v}_{i}$, and then feed them through the prediction layer to obtain the estimated distribution:
$q_{i,l}=g_{\text{PLM}}([\hat{w}_{i,l}\oplus\textbf{v}_{i}])\in\mathbb{R}^{V},$ (5)
where $\oplus$ is the concatenation operation, $V$ is the vocabulary size, and $g_{\text{PLM}}$ is a two-layer MLP similar to the decoder of language model. Finally, similar to common masked language modeling [19], we leverage the cross-entropy loss for pretraining optimization:
$\mathcal{L}^{MLM}_{i}=\frac{1}{|\mathcal{I}^{text}|}\sum\nolimits_{l\in\mathcal{I}^{text}}\operatorname{CrossEntropy}\left(q_{i,l},\ x^{text}_{i,l}\right),$ (6)
where $x^{text}_{i,l}$ is the original $l$-th token.
IV-C4 Joint Reconstruction for Masked Tabular Modeling
Likewise, the MTM module takes the text-tabular pair $(\textbf{w}_{i},\hat{\textbf{v}}_{i})$ as input, and aims to recover the masked features. To dynamically capture the essential knowledge from the corresponding tokens of the masked features, we design the cross-attention unit to aggregate the tabular representation $\hat{\textbf{v}}_{i}$ with the “reference answer” $\textbf{w}_{i}$:
$\textbf{u}_{i}=\text{Softmax}\left(\frac{\widehat{\textbf{v}}_{i}\textbf{Q}\textbf{w}_{i}^{T}}{\sqrt{D_{text}}}\right)\textbf{w}_{i},\;\;\textbf{u}_{i}\in\mathbb{R}^{L}$ (7)
where $\textbf{Q}\in\mathbb{R}^{D_{tab}\times D_{text}}$ is the trainable cross-modal attention matrix, and $\sqrt{D_{text}}$ is the scaling factor [73].
For each masked feature with index $f\in\mathcal{I}^{tab}$, we maintain an independent MLP network $g_{\text{ID}}^{(f)}$ followed by a softmax function to compute the distribution $p_{i,f}\in\mathbb{R}^{M}$ over the candidate features:
$\textbf{c}_{i,f}$ $=g_{\text{ID}}^{(f)}(\textbf{u}_{i}),\quad\textbf{c}_{i,f}\in\mathbb{R}^{M},$ (8)
$p_{i,f,j}$ $=\frac{\text{exp}(c_{i,f,j})}{\sum_{k=1}^{M}\text{exp}(c_{i,f,k})},\quad j=1,\cdots,M,$
where $M$ is the size of the entire feature space. Finally, we adopt the cross-entropy loss on all the masked features:
$\mathcal{L}_{i}^{MTM}=\frac{1}{|\mathcal{I}^{tab}|}\sum\nolimits_{f\in\mathcal{I}^{tab}}\operatorname{CrossEntropy}(p_{i,f},\ x^{tab}_{i,f}),$ (9)
where $x^{tab}_{i,f}$ is the original feature of $f$-th field.
However, the loss above is actually impractical and inefficient since it has to calculate the softmax function over the entire feature space in Eq. 8, where $M$ is usually at million level for real-world applications. To this end, we adopt noise contrastive estimation (NCE) [74, 75, 7]. NCE transforms a multi-class classification task into a binary classification task, where the model is required to distinguish positive features (i.e., masked feature) from noise features. Specifically, for each masked feature $x_{i,f}^{tab}$ of $f$-th field, we sample $K$ noise features from the entire feature space according to their frequency distribution in the training set. Then, we utilize the binary cross-entropy (BCE) loss for MTM optmization:
$\mathcal{L}_{i}^{MTM}=-\frac{1}{|\mathcal{I}^{tab}|}\sum_{f\in\mathcal{I}^{tab}}(\log\sigma(c_{i,f,t})+\sum_{k=1}^{K}\log(1-\sigma(c_{i,f,k}))))$ (10)
where $\sigma$ is the sigmoid function, and $t$, $k$ are the indices of the positive and noise features respectively.
IV-C5 Instance-level Contrastive Learning (ICL)
In addition to the feature-level alignment brought by the joint reconstruction, we further introduce contrastive learning to explicitly establish instance-level alignment between two modalities. The contrastive objective draws the representations of matched text-tabular pairs together and pushes apart those non-matched pairs [34, 76].
In the ICL module, we leverage the [CLS] token vector $w_{i,1}$ to represent the textual input $\textbf{x}_{i}^{text}$. For dimensional consistency, we first employ two separate linear layers followed by a normalization layer to project the [CLS] token representation $w_{i,1}$ and tabular representation $\textbf{v}_{i}$ into $d$-dimensional vectors, $z^{text}_{i}$ and $z^{tab}_{i}$, respectively. $d$ is set to 128 in this paper. Next, we adopt InfoNCE [77] to compute the ICL loss:
$\mathcal{L}_{text2tab}^{ICL}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{j})/\tau)}\right),$ (11)
$\mathcal{L}_{tab2text}^{ICL}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{j})/\tau)}\right),$ (12)
$\mathcal{L}^{ICL}$ $=\frac{1}{2}(\mathcal{L}_{text2tab}^{ICL}+\mathcal{L}_{tab2text}^{ICL}),$ (13)
where $B$ is the batch size, $\tau$ is the temperature hyperparameter, and the similarity function $\text{sim}(\cdot)$ is measured by dot product.
Finally, by putting the three objectives together, the overall loss for the modality alignment pretraining stage is:
$\mathcal{L}^{pretrain}=\frac{1}{B}\sum\nolimits_{i=1}^{B}\left(\mathcal{L}^{MLM}_{i}+\mathcal{L}^{MTM}_{i}\right)+\mathcal{L}^{ICL}.$ (14)
[…]
IV-D Supervised Finetuning
After the pretraining stage, the PLM and ID-based model have received better parameter initialization with fine-grained cross-modal alignments. As depicted in Figure 5(a), in this stage, FLIP finetunes the two models jointly on the downstream CTR prediction task with supervised click signal to achieve superior performance.
FLIP places a randomly initialized linear layer on the ID-based model, and another layer upon the PLM, so that these two model can output the estimated probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ respectively.
$\hat{y}^{\text{ID}}_{i}$ $=\sigma(\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})),$ (15)
$\hat{y}^{\text{PLM}}_{i}$ $=\sigma(\operatorname{Linear}_{\text{PLM}}(\textbf{w}_{i}))$
And the final click probability is estimated by a weighted sum of outputs from both models:
$\hat{y}_{i}=\sigma\left(\alpha\times\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})+(1-\alpha)\times\operatorname{Linear}_{\text{PLM}}(\textbf{w}_{i})\right),$ (16)
where $\alpha\in$ is a learnable parameter to balance the outcomes from two different modalities. To avoid performance collapse mentioned in previous works [78, 79], we apply BCE objectives over the jointly estimated click probability $\hat{y}$, as well as the solely estimated click probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ for model optimization:
$\mathcal{L}^{finetune}=\mathcal{L}_{BCE}(y,\hat{y})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{ID}})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{PLM}})$ (17)
[…]
$\mathcal{L}_{ID}$ $=\mathcal{L}_{BCE}(y,\hat{y}^{\text{ID}}).$ (18)
[…]
$\mathcal{L}_{PLM}$ $=\mathcal{L}_{BCE}(y,\hat{y}^{\text{PLM}}).$ (19)
[…]
V Experiment
In this section, we aim to evaluate the performance and effectiveness of the proposed method. Specifically, we conduct extensive experiments to answer the following research questions:
  • RQ1
    Does our proposed FLIP outperform existing state-of-the-art baseline models for CTR prediction?
  • RQ2
    Does FLIP have high compatibility with various ID-based models and PLMs?
  • RQ3
    How do the main components (e.g., MLM or MTM) of FLIP affect the performance?
  • RQ4
    How do different hyperparameter settings (e.g., the mask ratio) of FLIP impact the performance?
  • RQ5
    Does FLIP really establish the feature-level alignment?
  • RQ6
    Does FLIP promote the feature embedding learning, and alleviate the representation degeneration problem?
V-A Experiment Setup
V-A1 Datasets
We conduct extensive experiments on three real-world public datasets. All of the selected datasets contain user and item information. The information is unencrypted original text, thus preserving the real semantic information.
  • MovieLens-1M [80] is a movie recommendation dataset with user-movie ratings ranging from 1 to 5. Following the previous work [17, 25], We consider samples with ratings greater than 3 as positive, samples with ratings less than 3 as negative, and remove samples with ratings equal to 3 (i.e., neutral).
  • BookCrossing [81] is a book recommendation dataset and possesses user-book ratings ranging from 0 to 10. We consider samples with scores greater than 5 as positive, and the rest as negative.
  • GoodReads [82, 83] is a book recommendation dataset which contains user-book ratings ranging from 1 to 5. We take samples with ratings greater than 3 as positive, and the rest as negative.
The statistics of the processed datasets are shown in Table I. Following previous works [25, 84], we sort all samples in chronological order and take the first 90% samples as the training set and the remaining as the testing set. The training set for the pretraining stage and finetuning stage are the same [7]. All our experimental results are obtained on the testing set.
TABLE I: Statistics of processed datasets.
Dataset #Samples #Fields #Features
MovieLens-1M 739,012 8 16,849
BookCrossing 1,031,171 8 722,235
GoodReads 20,122,040 15 4,565,429
V-A2 Evaluation Metrics
We adopt two popular metrics to evaluate the performance of different methods on the CTR task, following [14, 17]. AUC (Area Under the ROC Curve) measures the probability that the CTR predictor will assign a higher score to a randomly chosen positive item than a randomly chosen negative item. A higher AUC indicates a better performance. Logloss is the binary cross-entropy loss. A smaller Logloss indicates a better performance. Notably, a slightly higher AUC or a lower Logloss (e.g., 0.001) can be considered as a significant improvement in CTR prediction [45, 5, 85].
V-A3 Baselines
The baseline methods can be mainly classified into three categories: (1) ID-based models, (2) PLM-based methods, and (3) methods that combine both ID-based models and PLMs.
(1) ID-based models that take tabular data as input:
  • AFM [18] extends FM by learning the weights of feature interactions via attentional networks.
  • PNN [13] is a product-based network that feeds the inner products of features embeddings into a DNN.
  • Wide&Deep [43] is a general learning framework that combines a wide network and a deep network to achieve the advantages of both.
  • DCN [44] improves Wide&Deep by replacing the wide part with a cross network to learn explicit high-order feature interactions.
  • DeepFM [5] is a extension of Wide&Deep by replacing the wide network with FM to explicitly learn second-order feature interactions.
  • xDeepFM [45] uses a compressed interaction network (CIN) to capture high-order feature interactions in a vector-wise way.
  • AutoInt [17] leverages self-attention networks to learn high-order features interactions.
  • DCNv2 [14] modifies DCN by proposing a more expressive cross network to better capture explicit feature interactions.
  • AFN [46] adopts a logarithmic transformation layer to learn adaptive order feature interactions.
(2) PLM-based methods that take textual data as input:
  • CTR-BERT [22] is a two-tower structure that feeds user information and item information into their respective BERT [19] to make CTR predictions.
  • P5 [23] transforms various recommendation tasks into text generation tasks by employing the pretrained language model T5 [58]. For a fair comparison, we use P5 for a single task only (i.e., CTR prediction) with multiple templates.
  • PTab [21] pretrains the BERT model on masked language modeling task and finetunes this BERT on downstream prediction tasks.
(3) Methods that combine ID-based models and PLMs:
  • CTRL [25] is currently a state-of-the-art method for combining ID-based models and PLMs through instance-level knowledge distillation.
TABLE II: The overall performance of different models on three datasets. The best result is given in bold, and the second-best value is underlined. We also use wavy underline to denote the best baseline performance. Rel.Impr denotes the relative AUC improvement rate of FLIP against each baseline. The symbol $\ast$” indicates statistically significant improvement of FLIP over the best baseline with $p$-value $<$ 0.001.
Model MovieLens-1M BookCrossing GoodReads
AUC Logloss Rel.Impr AUC Logloss Rel.Impr AUC Logloss Rel.Impr
AFM 0.8449 0.3950 2.04% 0.7946 0.5116 1.45% 0.7630 0.5160 2.06%
PNN 0.8546 0.3946 0.88% 0.7956 0.5131 1.32% 0.7725 0.5055 0.80%
Wide&Deep 0.8509 0.3957 1.32% 0.7951 0.5116 1.38% 0.7684 0.5090 1.34%
DCN 0.8509 0.4056 1.32% 0.7957 0.5108 1.31% 0.7693 0.5086 1.22%
DeepFM 0.8539 0.3905 0.96% 0.7947 0.5122 1.43% 0.7671 0.5138 1.51%
xDeepFM 0.8454 0.3934 1.98% 0.7953 0.5108 1.36% 0.7720 0.5079 0.87%
AFN 0.8525 0.3868 1.13% 0.7932 0.5139 1.63% 0.7654 0.5118 1.74%
AutoInt 0.8509 0.4013 1.32% 0.7953 0.5118 1.36% 0.7716 0.5071 0.92%
DCNv2 0.8548 0.3893 0.85% 0.7956 0.5103 1.32% 0.7724 0.5057 0.82%
CTR-BERT 0.8304 0.4131 3.82% 0.7795 0.5300 3.41% 0.7385 0.5316 5.44%
P5 0.8304 0.4173 3.82% 0.7801 0.5261 3.33% 0.7365 0.5336 5.73%
PTab 0.8426 0.4195 2.31% 0.7880 0.5384 2.30% 0.7456 0.5268 4.44%
CTRL 0.8572 0.3838 0.57% 0.7985 0.5101 0.95% 0.7741 0.5045 0.59%
[…]
FLIP (Ours) 0.8621* 0.3788* - 0.8061* 0.5004* - 0.7787* 0.5001* -
V-A4 Implementation Details
All of our experiments are performed on 8 NVIDIA Tesla V100 GPUs with PyTorch [86]. In the modality alignment pretraining stage, we set the text and tabular mask ratio $r_{text}$ and $r_{tab}$ both to 15%. Unless specified otherwise, we adopt TinyBERT [87] as the basic PLM, and DCNv2 [14] as the basic ID-based model. During pretraining, the model is trained for 30 epochs with the AdamW [88] optimizer and a batch size of 1024. The learning rate is initialized as 5e-5 followed by a cosine decay strategy. The number of noise features $K$ is 25 in Eq. 10. The temperature $\tau$ is 0.7 in Eq. 13. In the supervised finetuning stage, we adopt the Adam [89] optimizer with learning rate selected from {1e-5,5e-5,1e-4,5e-4,1e-3,5e-3}. The finetuning batch size is 256 for MovieLens-1M and BookCrossing, and is 2048 for GoodReads.
For all ID-based models in the baselines or FLIP framework, the embedding size is fixed to 32, and the size of DNN layers is [300,300,128]. The structure parameters and the size of PLMs in baselines are set according to their original papers. We also apply grid search to baseline methods for optimal performance.
V-B Performance Comparison (RQ1)
[…]
  • FLIP outperforms all baselines significantly and consistently. On MovieLens-1M, FLIP outperforms the best baseline by 0.005 AUC points. On BookCrossing, FLIP achieves up to 0.007 AUC improvements compared to the best baseline. On GoodReads, FLIP outperforms the best baseline by 0.004 AUC points. These results confirm the excellence of our proposed fine-grained alignment and jointly finetuning approach.
[…]
  • The baseline CTRL generally achieves the best performance among all the baselines. CTRL distills the PLM’s knowledge into an ID-based model and uses this ID-based model to make prediction, which ignores the PLM’s inherent capabilities, such as reasoning. Furthermore, it only focuses on instance-level alignment, resulting in learned representations with high similarity (Section V-G). Our FLIP combines ID-based models and PLMs through fine-grained alignment and model ensemble, effectively mitigating these limitations and improving performance.
V-C Compatibility Analysis (RQ2)
FLIP serves a model-agnostic framework that is compatible for various backbone models. Therefore, in this section, we investigate the model compatibility of FLIP in terms of both ID-based models and PLMs.
V-C1 Compatibility with ID-based models
[…]
  • By tuning both the ID-based model and PLM, FLIP achieves the best performance on different backbone ID-based models significantly and consistently, demonstrating the superior compatibility of FLIP in terms of ID-based models.
TABLE III: The compatibility analysis of FLIP with different ID-based models on MovieLens-1M. The PLM is fixed as TinyBERT. N/A means to train the raw ID-based model from scratch. For each ID-based model, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over the raw ID-based model with $p$-value $<$ 0.001.
ID-based Model Finetuning Strategy MovieLens-1M
AUC Logloss
DeepFM N/A 0.8539 0.3905
[…]
FLIP 0.8615* 0.3758*
AutoInt N/A 0.8509 0.4013
[…]
FLIP 0.8600* 0.3807*
DCNv2 N/A 0.8548 0.3893
[…]
FLIP 0.8621* 0.3788*
TABLE IV: The compatibility analysis of FLIP with different ID-based models on BookCrossing. The PLM is fixed as TinyBERT. N/A means to train the raw ID-based model from scratch. For each ID-based model, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over the raw ID-based model with $p$-value $<$ 0.001.
ID-based Model Finetuning Strategy BookCrossing
AUC Logloss
DeepFM N/A 0.7947 0.5122
[…]
FLIP 0.8033* 0.5031*
AutoInt N/A 0.7943 0.5118
[…]
FLIP 0.8011* 0.5050*
DCNv2 N/A 0.7956 0.5103
[…]
FLIP 0.8061* 0.5004*
V-C2 Compatibility with PLMs
Similarly, we fix the ID-based model as DCNv2, and train FLIP with three different PLMs, including TinyBERT with 14.5M parameters, RoBERTa-Base [20] with 125M parameters, and RoBERTa-Large [20] with 355M parameters. The results are summarized in Table V and  VI, from which we observe that:
[…]
  • Although a larger language model is expected to bring about more significant performance improvements, we find that such performance gains might gradually diminishes. Due to the page limitation and resource constraint, we leave the further analysis on such a phenomenon as future works (e.g., involving further larger PLMs).
TABLE V: The compatibility analysis of FLIP with different PLMs on MovieLens-1M. The ID-based model is fixed as DCNv2. For each type of PLM, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over raw DCNv2 model with $p$-value $<$ 0.001.
PLM Finetuning Strategy MovieLens-1M
AUC Logloss
DCNv2 0.8548 0.3893
[…]
FLIP 0.8621* 0.3788*
[…]
FLIP 0.8634* 0.3770*
[…]
FLIP 0.8650* 0.3757*
TABLE VI: The compatibility analysis of FLIP with different PLMs on BookCrossing. The ID-based model is fixed as DCNv2. For each type of PLM, the best result is in bold, and the second-best is underlined. $\ast$” indicates statistically significant improvement over raw DCNv2 model with $p$-value $<$ 0.001.
PLM Finetuning Strategy BookCrossing
AUC Logloss
DCNv2 0.7956 0.5103
[…]
FLIP 0.8061* 0.5004*
[…]
FLIP 0.8083* 0.4995*
[…]
FLIP 0.8092* 0.4986*
Figure 6: The hyperparameter study on textual mask ratio $r_{text}$ and tabular mask ratio $r_{tab}$ on MovieLens-1M (left two images) and BookCrossing (right two images).
V-D Ablation Study (RQ3)
We conduct comprehensive ablation experiments for better understanding the contributions of different components in our proposed FLIP. In this section, all the experiments are conducted on Movielens-1M and BookCrossing datasets.
Firstly, we evaluate the impact of pretraining objectives (i.e., MLM, MTM, ICL) by eliminating them from the pretraining stage. Note that removing all three objectives means that the pretraining stage does not exist. The results are reported in Table VII.
  • As we can observe, the optimal performance is achieved when three losses are deployed simultaneously, and removing each loss will degrade performance, while eliminating all losses results in the lowest performance. These phenomena demonstrate every component contributes to the final performance.
  • Removing ICL (w/o ICL) obtains better performance than removing MLM&MTM (w/o MLM&MTM), indicating that joint modeling for MLM&MTM tasks can learn and align meaningful representations even without ICL.
Next, to further investigate the effect of joint reconstruction for MLM and MTM, we design the following two variants:
  • w/o Field-level Masking: We replace the field-level masking for textual data with the common random token-level masking, as discussed in Section IV-C1.
  • w/o Joint Reconstruction: The reconstruction of one masked modality no longer relies on the help from the other modality.
The results are shown in Table VII. The performance drops, especially on the Logloss metric, when we either remove the field-level masking strategy or eliminate the joint reconstruction. Such a phenomenon demonstrates the importance of fine-grained feature-level alignment, which ensures the feature-level communication and interaction between ID-based models and PLMs for dual modalities.
TABLE VII: The results of ablation study. The best values are given in bold.
Model Variant MovieLens-1M BookCrossing
AUC Logloss AUC Logloss
FLIP 0.8621 0.3788 0.8061 0.5004
w/o MLM 0.8610 0.3791 0.8042 0.5035
w/o MTM 0.8615 0.3806 0.8053 0.5032
w/o ICL 0.8618 0.3790 0.8039 0.5026
w/o MLM&MTM 0.8598 0.3815 0.8020 0.5044
w/o MLM&MTM&ICL 0.8593 0.3810 0.8008 0.5061
w/o Field-level Masking 0.8605 0.3785 0.8054 0.5015
w/o Joint Reconstruction 0.8618 0.3820 0.8050 0.5036
Figure 7: The hyperparameter study on the temperature $\tau$.
(a) FLIPw/oMLM&MTM
(b) FLIPw/oMLM
(c) FLIPw/oMTM
(d) FLIP
Figure 8: The visualization of feature representations learned by different model variants on MovieLens-1M dataset. We use SVD to project the feature embedding matrix into 2D data. Note that FLIPw/oMLM&MTM indicates the instance-level contrastive learning (ICL) variant without the MLM and MTM pretraining objectives.
Figure 9: The visualization of similarities between the sample representations of masked textual and tabular data. “Text-$f$” and “Tab-$f$” denote that we mask the $f$-th field of the input data of textual or tabular modalities, respectively.
V-E Hyperparameter Study (RQ4)
In this section, we conduct experiments to study the influence of different hyperparameters on the FLIP.
V-E1 The Impact of Mask Ratio $r_{text}$ and $r_{tab}$
Since there are two independent mask ratio for textual and tabular modalities, we fix the mask ratio on the one side and alter the mask ratio on the other side from $\{0.15,0.3,0.6\}$. The results are in Figure 6, from which we observe that the best performance is generally achieved when both mask ratios are relatively small (i.e., 0.15). The reason is that excessive masking might lead to ambiguity in the target modality data, which hurts the model pretraining [19, 20].
V-E2 The Impact of Temperature Hyperparameter $\tau$
The temperature $\tau$ controls the sharpness of estimated distribution. We select $\tau$ from $\{0.05,0.1,0.3,0.5,0.7,1.0\}$, and report the results in Figure 7. When the temperature gradually grows, the performance increases first and then decreases, with the optimal temperature choice between $0.5$ and $0.7$. As suggested in previous works [90, 91], a too small temperature would only concentrate on the nearest sample pairs with top scores, and a too large temperature might lead to a flat distribution where all negative sample pairs receive almost the same amount of punishment. Both of them will hurt the final performance [90, 25]. Therefore, we choose $\tau=0.7$ in our approach.
V-F Analysis on Feature-level Alignment (RQ5)
We conduct case studies on Movielens-1M and BookCrossing datasets to further explore the fine-grained feature-level alignment established during pretraining. For a dual-modality input pair $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$, we first perform field-level data masking to mask out each field respectively, resulting in $F$ pairs of corrupted inputs $\{(\textbf{x}_{i,(f)}^{text},\textbf{x}_{i,(f)}^{tab})\}_{f=1}^{F}$, where $\textbf{x}_{i,(f)}^{text}$ and $\textbf{x}_{i,(f)}^{tab}$ denote that we solely mask the $f$-field of the textual or tabular data. Then we employ the PLM and ID-based model to encode them into $F$ pairs of normalized sample representations $\{(\textbf{z}_{i,(f)}^{text},\textbf{z}_{i,(f)}^{tab})\}_{f=1}^{F}$. Next, we compute the mutual similarity scores (measured by dot product) over each cross-modal representation pair, and visualize the heat map in Figure 9.
We can observe that the similarity score varies a lot for cross-modal input pairs with different masked fields, and the top-similar score is achieved for pairs with the same masked field (i.e., on the diagonal). This indicates that FLIP can perceive the changes among field-level features for both modalities, and further maintain a one-to-one feature-level correspondence between the two modalities.
V-G Quality of Feature Representations (RQ6)
In this section, we investigate whether the feature-level alignment improves the feature embedding learning and therefore alleviates the representation degeneration problem. Following previous work [36], we adopt SVD [39] decomposition to project the learned feature embeddings of the pretrained ID-based model into 2D data. In Figure 8, we visualize the feature representations learned by four model variants FLIPw/oMLM&MTM, FLIPw/oMLM, FLIPw/oMTM and FLIP. Note that FLIPw/oMLM&MTM indicates the instance-level contrastive learning (ICL) variant without the MLM and MTM pretraining objectives. We have the following observations:
  • The feature embeddings learned by FLIPw/oMLM&MTM (i.e., ICL) collapse into a narrow cone, suffering from severe representation degeneration problem. In contrast, FLIP obtains feature embeddings with more distributed latent patterns, thus better capturing the feature diversity. This highlights the effectiveness of feature-level alignment in promoting representation learning and mitigating the issue of representation degeneration.
  • Comparing FLIP with $\text{FLIP}_{w/oMLM}$ or $\text{FLIP}_{w/oMTM}$, we find that removing either MLM or MTM objective makes the learned embeddings more indistinguishable, demonstrating the necessity of dual alignments between modalities for learning effective representations.
VI Conclusion
In this paper, we propose FLIP, a model-agnostic framework that achieves the fine-grained alignment between ID-based models and PLMs. We view tabular data and transformed textual data as dual modalities, and design a novel modality alignment pretraining task. Specifically, the joint reconstruction for masked language/tabular modeling (with specially designed masking strategies) and cross-modal contrastive learning are employed to accomplish feature-level and instance-level alignments, respectively. Furthermore, we propose to jointly finetune the ID-based model and PLM to achieve superior performance by combining the advantages of both models. Extensive experiments show that FLIP outperforms state-of-the-art baselines on three real-world datasets, and is highly compatible with various ID-based models and PLMs.